OpenStudy (anonymous):

Evaluate cos^2 θ for cos θ = sqrt(3)/2

zepdrix (zepdrix):

\[\Large\rm \cos \theta=\frac{\sqrt3}{2}\]Soooooo, they want you to square it....? Is that all?\[\Large\rm \cos^2\theta=\left(\frac{\sqrt{3}}{2}\right)^2\]\[\Large\rm \cos^2\theta=\frac{3}{4}\]Or do we need to do something more with this? 0_o Confusing directions...

OpenStudy (anonymous):

Oh that was it thank you! I just didn't understand how to do it

zepdrix (zepdrix):

:3

zepdrix (zepdrix):

Do you understand why \(\Large\rm \sqrt{3}\) gives us 3 when we square it? :U

OpenStudy (anonymous):

Yeah

zepdrix (zepdrix):

k good c: